OpenStudy (anonymous):

REALLY need help with 3log3 (x+2)=9 or solve x^4+3x^3-2x^2+48x= 288 if one root is x+4i

OpenStudy (anonymous):

is that log base 3 or log^3?

OpenStudy (anonymous):

log base 3

OpenStudy (anonymous):

ok, for the other one if x+4i is a root, x-4i also is one, see http://www.wolframalpha.com/input/?i=factor+x^4%2B3x^3-2x^2%2B48x-+288+

OpenStudy (anonymous):

you could use polynom division twice for that one, for the one with the log simply convert to log_3((x+2)^3)=9

OpenStudy (anonymous):

then 3^ so (x+2)^3 =3^9, so x+2=3^3, solve for x

OpenStudy (anonymous):

thank you very much, I haven't done this stuff all summer and have a sheet on it all due tomorrow